OpenStudy (anonymous):

Given ΔABC with A(–3, 2), B(–1, –4), and C(4, 1), write the equation for the line containing altitude in standard form. (1 point)x – y = 1 x + y = –1 x + y = 1 x – y = –1

OpenStudy (anonymous):

could someone pleaae help me on this question

OpenStudy (anonymous):

PLLLLLLLEEEESSSEE HELP ME!!!!!!!!!!!!!!!!!!!!

OpenStudy (mathstudent55):

 There are 3 different altitudes. Does the problem ask you for which one?

OpenStudy (anonymous):

no it just says to write the equation for the line containing altitude in standard form.

OpenStudy (anonymous):

. Given ΔABC with A(–3, 4), B(5, 8), and C(2, –2), write the equation of the line containing midsegment in standard form, where X is the midpoint of and Z is the midpoint of . (1 point)6x – 5y = –24 2x + 5y = 32 2x – 5y = –28 6x + 5y = 36

OpenStudy (anonymous):

could you help me on that one to pllleeaasse

OpenStudy (anonymous):

mathstudent55??????? mathstudent55??????

OpenStudy (mathstudent55):

What are X and Z the midpoints of? You copy problems and paste them here, but the characters that don't get copied, you just leave out. There's information missing here too.

OpenStudy (anonymous):

for the first one AR is in between "altitude" and "in standard form"

OpenStudy (mathstudent55):

That's better.